### v1

# Foundation Models for Autonomous Driving System: An Initial Roadmap

Abstract.
Recent advancements in Foundation Models (FMs), such as Large Language Models (LLMs), have significantly enhanced Autonomous Driving Systems (ADSs) by improving perception, reasoning, and decision-making in dynamic and uncertain environments. However, ADSs are highly complex cyber-physical systems that demand rigorous software engineering practices to ensure reliability and safety. Integrating FMs into ADSs introduces new challenges in system design and evaluation, requiring a systematic review to establish a clear research roadmap. To unlock these challenges, we present a structured roadmap for integrating FMs into autonomous driving, covering three key aspects: the infrastructure of FMs, their application in autonomous driving systems, and their current applications in practice. For each aspect, we review the current research progress, identify existing challenges, and highlight research gaps that need to be addressed by the community.
Foundation Model, Autonomous Driving System, Roadmap, V2X
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYbooktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwayisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Autonomous driving systems face increasingly complex challenges in navigating diverse real-world scenarios. Traditional approaches, while effective in controlled environments, often struggle with unseen situations, unexpected obstacles, and dynamic interactions that characterize real-world driving conditions (Wu et al., 2024). These limitations stem primarily from their reliance on predetermined rules and supervised learning from finite labeled datasets, which cannot fully encompass the spectrum of scenarios a vehicle might encounter in operation.
The emergence of foundation models (FMs) trained on vast and diverse datasets has demonstrated unprecedented capabilities in reasoning and generalization across various domains (Bommasani et al., 2022). These models have exhibited remarkable abilities in understanding context, reasoning about the context, and generating appropriate responses. Their success in natural language processing (Devlin et al., 2019; Touvron et al., 2023) and computer vision (Radford et al., 2021; Liu et al., 2023b; OpenAI, 2023) tasks suggests promising applications in addressing the fundamental challenges of autonomous driving.
The autonomous driving ecosystem stands to benefit significantly from the integration of FMs, which can enhance real-world scenario understanding, improve decision-making, and facilitate robust system development. For instance, FMs can leverage their broad knowledge base to interpret complex traffic scenarios, understand natural language instructions from passengers, and make informed decisions in previously unseen situations (Zhou et al., 2024b). This integration could bridge the gap between traditional autonomous driving systems’ capabilities and the requirements for truly robust autonomous operation in diverse real-world conditions.
To unlock these benefits, we propose a structured roadmap for integrating FMs into autonomous driving, covering three key aspects: FM infrastructure, their integration across autonomous driving system modules, and their practical real-world applications, as shown in Figure 1. For each aspect, we review the current research progress, identify existing challenges, and highlight research gaps that need to be addressed by the community. Through this comprehensive analysis, we aim to guide future research efforts in leveraging FMs to advance the field of autonomous driving.
Figure 1. Overview of the roadmap.
[…]
Topics covered in this paper.
2. FM4AD Infrastructure
The infrastructure is the cornerstone for integrating FMs into autonomous driving, encompassing the datasets, computational resources, and toolchains necessary for training, testing, and deployment of FMs.
2.1. High-quality Dataset for Autonomous Driving
High-quality datasets have played a critical role in advancing autonomous driving technology. Traditional datasets primarily focused on 2D annotations like bounding boxes and masks for RGB camera images (Pan et al., 2018). With the emergence of FMs, datasets are evolving towards multi-modal integration, particularly incorporating language descriptions (Zhou et al., 2024b). While this multi-modal approach promises to accelerate autonomous driving development, it introduces new challenges. Moreover, the massive data requirements for training FMs raise significant concerns about privacy protection and ethical/legal compliance (Longpre et al., 2024). Thus, we identify the following challenges:
Challenge I: Dataset Cleaning and Curation. Dataset cleaning and curation serve as a critical foundation for developing FMs, ensuring data integrity, privacy protection, and efficient training. Key challenges include protecting privacy (Zhu et al., 2024a; Zhang et al., 2024b; Carlini et al., 2021), and mitigating bias in training datasets (Llorca et al., 2024; Li et al., 2024a; Hort et al., 2024a). The privacy challenge involves both preventing personal data from appearing in training datasets and ensuring it cannot be inferred from model outputs (Kim et al., 2023; Zhu et al., 2024a). Bias in training data can lead models to perpetuate systemic inequities and may introduce safety risks for unrepresented groups when deployed in real-world scenarios (Llorca et al., 2024). These challenges present several key opportunities for further research.
  • Opportunity: Bias Mitigation. Dataset bias poses a critical challenge for autonomous driving. Recent studies (Llorca et al., 2024) have revealed limited diversity in key demographic attributes (i.e., age, sex, and skin tone) within existing AD datasets. This lack of representation could lead to safety risks when deploying FMs trained on such datasets, particularly for underrepresented groups. While recent research has made progress in addressing these concerns (Li et al., 2024a; Hort et al., 2024a, b; Chen et al., 2025), the detection and mitigation of dataset bias remains an important open challenge in autonomous driving.
  • Opportunity: Privacy Preservation. Another critical research opportunity lies in developing effective privacy-preserving algorithms for FMs in autonomous driving. Established techniques such as differential privacy (Duan et al., 2023; Chua et al., 2024; Tang et al., 2024a; Tramèr et al., 2024), data cleaning (Kandpal et al., 2022; Carlini et al., 2021), and federated learning (Xu et al., 2024a; Zhang et al., 2024b; Sun et al., 2024b), have advanced the field, yet they consistently face challenges in balancing privacy preservation with data utility. The challenge is particularly acute with FMs, which tend to memorize training data extensively, potentially leading to privacy leakage even with the data used in fine-tuning processes (Zhu et al., 2024a). There is an urgent need for novel techniques that can ensure robust privacy guarantees while maintaining the comprehensive nature of training datasets required for large FMs.
Challenge II: Augmenting Autonomous Driving Datasets. Despite significant investments in the development of autonomous driving datasets, current limitations in data quality and scale hinder their ability to comprehensively address the field’s challenges (Cui et al., 2024c). Moreover, certain critical scenarios remain difficult or nearly impossible to capture in real-world data collection (Wei et al., 2024c; Song et al., 2024a). These include high-risk situations such as accident aftermath and pedestrian-involved incidents. However, comprehensive testing of autonomous vehicles against these scenarios is crucial for safety validation. To overcome these challenges, researchers should increasingly explore methods for generating customized driving scenarios, enabling the effective simulation of these critical cases to augment existing datasets and enhance their utility.
  • Opportunity: Customizable Driving Scenario Generation. Current autonomous driving systems are primarily trained and evaluated on datasets collected from daily driving scenarios or synthetic data (Ding et al., 2023; Wei et al., 2024c). However, these datasets generally lack safety-critical scenarios that are crucial for robust system evaluation. Research in driving scenario generation has progressed along multiple directions, including data-driven approaches (Kruber et al., 2019; Tan et al., 2021; Pronovost et al., 2023), adversarial generation methods (Ding et al., 2020; Wang et al., 2021; Abeysirigoonawardena et al., 2019; Mei et al., 2025), and knowledge-based techniques (Zhang et al., 2024c; Tang et al., 2024b; Wei et al., 2024c). Looking ahead, scenario generation algorithms need to address key challenges, such as maintaining consistency across multiple sensor modalities (e.g., LiDAR, camera images) and enhancing scenario complexity through interaction and collaboration between agents.
Challenge III: Dataset Licensing and Management. Dataset licensing and management pose a variety of challenges vital to ensuring the legal and ethical use of autonomous driving datasets. The massive amount of data required for training FMs heightens the risks of copyright breaches, licensing violations, and subsequent legal liabilities. Additionally, the terms of use for datasets released by leading autonomous driving companies vary widely, further complicating this task. The multimodal nature and diverse sources of autonomous driving datasets intensify these difficulties. Moreover, selecting/sampling the right training data is essential for producing capable FMs (Wettig et al., 2024; Brown et al., 2020; Chowdhery et al., 2023). Recent studies (Wolter et al., 2023; Kim et al., 2024b) have revealed the complex landscape of modern large dataset copyright and licensing, emphasizing the need for deeper exploration and development of innovative techniques. These challenges also open up opportunities for further research.
  • Opportunity: Dataset License Compliance. The primary challenge of license management lies in the complexity and variety of licenses governing autonomous driving datasets (Jahanshahi and Mockus, 2025; Xu et al., 2025). Unlike traditional datasets for LLMs, which primarily consist of publicly available data (crawled from the Internet) supplemented with proprietary datasets having usage restrictions, most autonomous driving datasets are released by leading autonomous driving companies with their own specific terms of use, necessitating careful review and understanding to ensure compliance (Wolter et al., 2023; Lucchi, 2024; Vendome et al., 2017). As pointed out by Kim et al. (Kim et al., 2024b), the scale of modern datasets renders manual compliance verification impractical, thereby requiring automated detection techniques. Promising research directions include the development of automated detection and audit systems for legal terms of use, providing developers with clear insights into the permissions and restrictions associated with each dataset.
  • Opportunity: Data Management Framework for FMs. As FMs demonstrate performance improvements through data scaling and the significance of data becomes evident (Chowdhery et al., 2023), effective data management becomes increasingly critical. While various tools and methods have been proposed to explore how to properly manage the training data, encompassing data deduplication (Lee et al., 2022a), training data selection (Rae et al., 2022; Wang et al., 2024a; Li et al., 2024b), sampling high-quality data (Peng et al., 2025; Wettig et al., 2024), and dataset license compliance (Kim et al., 2024b; Jahanshahi and Mockus, 2025), there is still a lack of a unified framework and criteria. Although there have been some initial attempts in this area (Wang et al., 2024c; Ostendorff et al., 2024; Peng et al., 2025), systematic approaches to data management for FMs remain in their early stages. Given the massive scale and diverse sources of data required for training FMs, developing a comprehensive data management framework has become an urgent priority.
2.2. Computational Resources
Due to the computational-intensive nature of FMs, computational resources, including graphic processing units (GPUs), tensor processing units (TPUs), and other specialized AI accelerators, form the very foundation of the FM infrastructure. Building upon this hardware layer, distributed training frameworks and cloud computing enable efficient resource utilization and management. However, the complexity of distributed, computation-intensive training, and reliable efficient deployment introduces unique challenges and opportunities in adopting FMs in autonomous driving. In this section, we mainly discuss the challenges and opportunities related to hardware, issues with the software layer (e.g., distributed training framework) are discussed in Section 2.3.
Challenge & Opportunity: Trustworthy Hardware Design. The computational-intensive nature of FM training and inference necessitates reliance on proprietary and specialized hardware solutions. However, this dependency complicates ensuring compatibility, optimization, and security across the hardware stack (Wang et al., 2024b). The proprietary nature of hardware design and firmware creates a ”black-box” environment, where potential vulnerabilities may remain undetected (Yang et al., 2016; Wang et al., 2024b). Furthermore, the parallel processing architecture and shared resources make GPUs particularly vulnerable to hardware attacks (Jiang et al., 2017), potentially resulting in sensitive information leakage (e.g., model parameters (Naghibijouybari et al., 2021)) or even enabling arbitrary code execution (Lee et al., 2022b). Given that security is crucial for autonomous driving systems, there is an urgent need to design secure hardware that incorporates security features at the hardware level and can effectively resist hardware-level threats such as side-channel attacks.
2.3. Development and Deployment of FM4AD
Due to their formidable size and computational demands, developing and deploying FMs has posed significant new challenges for autonomous driving applications. In this section, we discuss the challenges and opportunities related to the development and deployment of FMs for autonomous driving.
Challenge & Opportunity: Understanding FM Development Toolchain. The development toolchain for FMs presents unprecedented complexity compared to traditional deep learning frameworks. The enormous scale of these models amplifies the intricacy of data pre-processing pipelines and distributed training systems (Wang et al., 2024b). Additionally, the fast-evolving nature of FM development hinders developers and researchers from maintaining a comprehensive understanding of the continuously evolving toolchain. A promising research avenue lies in mining the FM development toolchain. This approach entails a thorough examination and assessment of the libraries and tools employed in FM development. By analyzing the toolchain, developers can uncover inefficiencies and streamline processes while preserving functionality (Wang et al., 2024b; Nguyen et al., 2019; Liu et al., 2023a). Furthermore, this exploration may reveal gaps, fostering the design of innovative tools or enhancements to existing ones, better addressing the growing demands of FM development.
Challenge & Opportunity: Efficient and Reliable FM Deployment in Vehicle. With growing concerns regarding privacy and strict response time requirements for autonomous driving tasks, efficiently deploying FMs in vehicles has become increasingly important. Existing approaches have investigated model compression techniques such as pruning (Ma et al., 2023; Sun et al., 2024a), knowledge distillation (Gu et al., 2024; Kim et al., 2024a), and model quantization (Lin et al., 2024; Xi et al., 2025; Li et al., 2025), alongside inference optimization techniques including parallel computation (Shoeybi et al., 2020; Yu et al., 2024), KV cache (Kwon et al., 2023; Luohe et al., 2024), and request scheduling (Han et al., 2022; Yang et al., 2024b). While these techniques can alleviate computational burden and improve response times, they may introduce new vulnerabilities. For instance, researchers have identified tailored attacks targeting quantized models (Zhang et al., 2025; Yang et al., 2024a) and KV cache-based optimizations (Song et al., 2025; Wu et al., 2025). Consequently, there is an urgent need to develop techniques that ensure both safety and efficiency when deploying FMs in vehicles.
Challenge & Opportunity: Edge/Cloud Collaboration for FM Services. While quantization techniques can alleviate the computational burden of edge devices, model capability remains limited by available edge computing power. Conversely, cloud computing infrastructure offers high-performance processing for complex tasks but often struggles to meet privacy, reliability, and latency requirements crucial for autonomous driving. Recent research efforts have been focused on investigating collaboration between edge and cloud (Yang et al., 2024c; Hao et al., 2024; Zhang et al., 2024a; Tian et al., 2024b). By developing frameworks that intelligently schedule and coordinate tasks between edge and cloud resources, autonomous driving systems can achieve collaborative intelligence, enhance adaptability to varying conditions, leverage edge knowledge while preserving privacy, and optimize resource utilization across the edge-cloud continuum. Potential research directions include optimizing heterogeneous architecture fusion strategies, designing asynchronous update mechanisms, and robust communication schemes.
3. Foundation Models in Vehicle
In this section, we examine how FMs enhance different modules of autonomous driving, summarizing techniques and methodological advances. Specifically, we mainly focus on how FMs can help achieve human-like driving using LLMs, VLMs, and world model-based prediction. We also identify key challenges and research opportunities to guide future investigations in this rapidly evolving field.
3.1. FM-Enhanced Autonomous Driving
Existing integration of FMs into autonomous driving systems can be roughly categorized into perception and scene understanding, navigation and planning, decision-making and control, and end-to-end autonomous driving. In this section, we briefly discuss representative techniques and identify the challenges and opportunities. For more background and technique details, we refer readers to prior works (Wu et al., 2024; Gao et al., 2024; Cui et al., 2024c; Zhou et al., 2024b).
3.1.1. Perception and Scene Understanding
FMs enhance perception in autonomous driving by enabling context-aware environmental understanding (Zhou et al., 2024b). VLMs such as LLaVA (Liu et al., 2023b) and GPT-4V (OpenAI, 2023) support tasks like 3D open-vocabulary object detection (Najibi et al., 2023; Pan et al., 2024b), language-guided retrieval (Romero et al., 2023; Wei et al., 2024a), and visual question answering (VQA) (Qian et al., 2024; Choudhary et al., 2024; Nie et al., 2025; Liu et al., 2024). Examples include OpenScene (Peng et al., 2023) for zero-shot 3D semantic segmentation and NuScenes-QA (Qian et al., 2024) for VQA benchmarks. Additionally, DriveVLM (Tian et al., 2024a) employs chain-of-thought reasoning for scene analysis, while DriveDreamer (Wang et al., 2025) predicts future states for proactive responses.
Challenge & Opportunity: Hallucination Mitigation. While FMs (e.g., LLMs and VLMs) have achieved significant advancements in autonomous driving tasks, hallucination remains a critical challenge for real-world deployment. FMs are particularly prone to hallucination, which refers to generating outputs that are factually incorrect, inconsistent, or nonsensical (Sun et al., 2025; Chakraborty et al., 2025). For instance, in the autonomous driving context, hallucinated object detection, such as mistakenly identifying a non-existent pedestrian, could lead to severe safety incidents like abrupt stops and potential collisions. Although substantial research has addressed hallucination in LLMs (Song et al., 2024b; Yao et al., 2024; Huang et al., 2025, 2024; Zhou et al., 2024a; Arteaga et al., 2024) and in FM for autonomous driving (Dona et al., 2024; Fan et al., 2024), the underlying triggers and effective detection methods remain unclear (Xu et al., 2023). Potential research directions include leveraging multi-modal FMs to ground language with visual information (Chakraborty et al., 2025) and developing mitigation strategies that preserve the models’ generation and reasoning capabilities.
3.1.2. Navigation and Planning
FMs integrate natural language into navigation and planning by converting textual instructions into spatial representations. Systems like Talk to the Vehicle (Sriram et al., 2019) and Ground then Navigate (Jain et al., 2023) generate waypoints and trajectories from multi-modal inputs. ALT-Pilot (Omama et al., 2023) enhances planning with language-augmented maps using CLIP (Radford et al., 2021), while GPT-Driver (Mao et al., 2023) and DriveVLM (Tian et al., 2024a) support predictive planning and reasoning.
3.1.3. Decision-Making and Control
FMs improve decision-making and control by translating scene understanding into safe actions. LLMs in Drive as You Speak (Cui et al., 2024a) and LanguageMPC (Sha et al., 2023) process complex data for real-time decisions. Hybrid systems like BEVGPT (Wang et al., 2024d) and Driving with LLM (Chen et al., 2024) combine reasoning with traditional controls, while SurrealDriver (Jin et al., 2024b) and Drive Like a Human (Fu et al., 2024) enhance robustness through safety and memory modules.
3.1.4. End-to-End Autonomous Driving
Recent advancements in FMs have enabled the development of unified models that integrate perception, reasoning, and control into a single differentiable framework. DriveGPT4 (Xu et al., 2024b) processes sensor inputs and queries for control signals and explanations. ADAPT (Jin et al., 2023) maps video to actions and narratives, DriveMLM (Wang et al., 2023) integrates LLMs into closed-loop systems, and VLP (Pan et al., 2024a) promotes generalization with context-aware frameworks.
Challenge & Opportunity: Multi-modality Adaptation. Foundation models, particularly LLMs, and MLLMs, have demonstrated remarkable reasoning capabilities in autonomous driving tasks. However, most existing approaches heavily rely on environmental information from upstream perception modules, making them vulnerable to input errors (Gao et al., 2024). Even minor perception inaccuracies, such as inaccurate object heading estimation, can lead to catastrophic failures in decision-making (Mao et al., 2024). This highlights the critical need for research into robust adaptation methods that can better handle uncertainties and errors in perception inputs while maintaining reliable decision-making capabilities.
Challenge & Opportunity: Domain-Specific Foundation Models for Autonomous Driving. While the open-source landscape for code-centric large language models (LLMs) has thrived with examples like Magicoder (Wei et al., 2024b) and CodeLlama (Rozière et al., 2024) setting benchmarks, most FMs for autonomous driving remain proprietary, such as GAIA-1 (Hu et al., 2023). This restricts academic and independent researchers from advancing innovation in a field where safety and robustness are critical. The core problem is the absence of a pre-trained, powerful foundation model designed for the specific tasks of autonomous driving, such as integrating multi-modal data (e.g., cameras, LIDAR) and handling complex decision-making in dynamic environments. An open-source, domain-specific foundation model is urgently needed to bridge this gap. By providing a robust starting point for tasks like perception, planning, and control, this model would empower researchers to address real-world driving challenges efficiently.
3.2. FM-Enabled Intelligent User Experience
Beyond enhancing core autonomous driving capabilities, FMs are revolutionizing user interaction and experience in autonomous vehicles. This subsection explores two key aspects: intelligent user interfaces with personalization, and enhanced surrounding awareness capabilities.
Challenge & Opportunity: Intelligent User Interface and Personalization. While FMs enable more intelligent and personalized user experiences in autonomous vehicles, several challenges need to be addressed. MLLMs like GPT-4V can interpret natural language instructions to control vehicles according to user preferences. For example, Cui et al. demonstrated that LLM-based planners can respond to personalized commands such as “drive aggressively,” adjusting vehicle behavior across different speeds and risk levels (Cui et al., 2024b). However, this flexibility raises significant safety concerns. As shown in (Cui et al., 2024d), LLMs may interpret and execute potentially dangerous commands like “drive as fast as you can.” Although research has explored methods to ensure compliance with traffic rules and safety requirements (Cui et al., 2024d; Yang et al., 2024c), the vulnerability to jailbreak attacks remains a concern, particularly given the proliferation of LLM exploitation techniques (Rao et al., 2024; Zou et al., 2023; Zhu et al., 2024b; Jin et al., 2024a). Additionally, balancing real-time responsiveness with user privacy presents another significant challenge, as discussed in Section 2.3.
Challenge & Opportunity: FM-Enabled Surrounding Awareness. FMs could enhance surrounding awareness by providing users with real-time, interpretable insights about the vehicle’s environment. For instance, DriveGPT4 (Xu et al., 2024b) integrates this awareness into the driving loop, offering passengers explanations for vehicle actions (e.g., “veers left to avoid collision”). This awareness extends to both safety and convenience features, such as alerting users to nearby hazards or points of interest, enhancing the overall experience (Domova et al., 2024; Miller et al., 2014). However, ensuring the accuracy and reliability of FM-generated insights remains challenging, as hallucinations or misinterpretations could mislead users. Additionally, presenting complex information requires careful UI design to maintain user-friendliness. Potential opportunities include developing robust multi-modal grounding techniques to reduce errors through cross-validation of visual and textual data, and creating intuitive visualization methods such as augmented reality overlays to effectively convey FM insights. These advancements could transform vehicles into intelligent companions that enhance both safety and user engagement.
4. Foundation Model Application in Practice
This section explores the practical deployment of FMs in autonomous driving. We distinguish between modular integration and full adoption of FMs, showcasing their role in enhancing vehicle capabilities.
Several initiatives employ FMs as specialized components within autonomous driving systems. Xiaomi SU7 integrates a VLM via OTA update to enhance scene interpretation and safety alerts (su7, [n. d.]). Li Auto combines a VLM with an end-to-end framework in its OTA-updated smart driving system, improving scene recognition and maneuver accuracy (lia, [n. d.]). TIER IV utilizes an LLM to enable vehicles to reason and communicate, enhancing human-vehicle interaction (tie, [n. d.]). Similarly, Bosch researchers apply natural language processing to predict traffic behaviors, boosting situational awareness (Keysan et al., 2023). These cases demonstrate FMs augmenting specific functions like perception and communication.
Meanwhile, full adoption leverages FMs as the core of autonomous driving systems. Cui et al. deploy an LLM in Talk2Drive for end-to-end control, personalizing driving through language and vision inputs (Cui et al., 2024d). Their subsequent work fully integrates a VLM onboard for motion control, unifying perception and decision-making (Cui et al., 2024e). Waymo’s MotionLM uses FMs to transform multi-agent motion prediction into a language task, streamlining dynamic interactions (Seff et al., 2023). These efforts highlight FMs driving comprehensive, adaptive autonomy.
Challenge: Foundation Model Alignment. As FMs become increasingly integrated into autonomous driving systems, their potential societal risks demand careful consideration. The undesired behaviors exhibited by FMs, such as hallucination, raise particular concerns in safety-critical domains like autonomous driving where they directly impact public safety. AI alignment has emerged as a potential solution, aiming to ensure AI systems behave in accordance with human intentions and values (Leike et al., 2018). Despite its critical importance for the safe deployment of FMs in autonomous driving, research in this area remains limited (Kong et al., 2024; IEEE, 2025; Abbo et al., 2024; Wolf et al., 2024). The complexity of foundation model systems, encompassing fairness, privacy, and security concerns, urgently calls for more attention and investigation into alignment strategies. Current alignment research can basically be divided into two key components: forward alignment and backward alignment (Ji et al., 2024), below are potential opportunities:
  • Opportunity: Enhancing Feedback Mechanisms (Forward Alignment). Forward alignment, which focuses on proactively shaping model behavior during training, presents a significant opportunity for improving FMs in autonomous driving. By incorporating human-value feedback during the training process, developers can construct more robust systems where FMs not only continuously learn but also maintain alignment with human intentions and safety requirements (Wang et al., 2024b).
  • Opportunity: Safety Benchmarks and Evaluation for Assurance (Backward Alignment). Datasets and benchmarks are crucial for safety evaluation, serving as fundamental tools for ensuring AI alignment. A key opportunity lies in developing comprehensive metrics and benchmarks for FMs to better evaluate their safety performance and ability to minimize accidents during task execution (Ji et al., 2024). Unlike traditional deep learning models, FMs can leverage general knowledge rather than actual cues to achieve unexpectedly high scores on existing metrics (Xie et al., 2025). This limitation highlights the need for more comprehensive benchmarks and metrics that can accurately assess both FM capabilities and potential deviations from intended behaviors (Amodei et al., 2016).
5. Conclusion
In this paper, we conducted a systematic exploration of integrating foundation models in autonomous driving, examining three key aspects: FM infrastructure, FM in Vehicles, and their real-world applications. For each aspect, we identified critical challenges and highlighted promising research opportunities. In conclusion, we believe that although significant challenges remain in integrating FMs into autonomous driving systems, it has shown its substantial potential for advancing the field. We hope this paper can serve as a roadmap for future research directions and accelerate the development of more capable autonomous driving systems.